# Exploring Citation Diversity in Scholarly Literature: An Entropy-Based Approach

### v1

@@ IV.3.2 Nobel Prize winners in Chemistry
From 1901 to 2023 the Nobel prize in chemistry has been bestowed on a total of 192 individuals. For our analysis, we have picked up the citation data of 30 Nobel prize winners in chemistry, 15 during 1901-1920 and 15 between 2018-2023. Table 8 represents the calculated diversity values and average citation counts of all the listed Nobel prize winners in chemistry. In Fig. 11, we see the citation diversity of 15 recent Nobel laureates in chemistry, with diversity values ranging from 60% to 90%. Two distinct groups emerge: one containing four laureates with citation diversity between 60%-70%, indicating higher citation inequality, and another between 80%-90%, reflecting a more balanced citation distribution. The minimal confidence intervals confirm the reliability of these values. Fig. 11(b) further shows that despite similar average citation counts, diversity values differ significantly, revealing more insightful patterns in citation distribution. Meanwhile, Fig. 12 shows earlier laureates’ citation diversity ranging from 65% to 90%, though with larger confidence intervals, indicating greater variability.
@@ IV.3.3 Abel prize winners in Mathematics
The Abel prize is awarded annually (2003 onwards) to one or more outstanding mathematicians and is widely considered the Nobel prize of mathematics. Here we consider the data for 15 Able prize winners during 2012-2023. Table 9 presents the average citation counts and diversity values for each Abel prize winner considered. In Fig. 13, we observe the citation diversity and box-plot for the citation counts of each Abel prize winner. These citation diversity values of each scientist range from 70% to 90%, indicating moderate citation inequality. The confidence intervals for most diversity values are not high, though a few have slightly higher confidence intervals. This suggests that most calculated diversity values are reliable, with some variability due to the use of publicly available data.
@@ IV.3.4 Turing Award winners in Computer Science
In Fig. 14, we present an overview of the diversity values and citation ranges for each Turing award winner in computer science. Notably, the diversity value of one of the 2015 Turing award winners is significantly lower compared to the others, indicating a high citation inequality for that individual. Conversely, the diversity values of the other computer scientists range between 55% and 85%, with a distinct division at 70%. Below this threshold, there are 7 Turing award winners, and above it, there are also 7 winners. This suggests that the citation inequality is higher for the 7 scientists below 70% compared to those above it. In this analysis, the confidence interval is minimal, except for two cases where it is slightly larger to increase the possibility range of that value. For better understanding, we maintain the same order of the award winners on the x-axis in Fig. 14(b) as in Fig. 14(a), where they are arranged in ascending order of diversity values. However, no meaningful pattern is observed from the total citation range in Fig. 14(b) but there is a clear pattern in diversity values of each scientists shown in Fig. 14(a).
@@ IV.3.5 Nobel prize winners in Economics
The first Nobel Memorial Prize in economic sciences was awarded in 1969. As of the awarding of the 2023 prize, 55 prizes in economic sciences have been given to 93 individuals. We considered the publication data of 15 distinguished economists between 2017-2023. Table 11 presents the diversity values and average citation counts for these 15 individual Nobel laureates in economics.
@@
Fig. 15(a) displays the citation diversity values for each laureate, showing that two recent laureates have very low diversity, indicating high citation inequality. In contrast, citation diversity of the other laureates ranging from 70% to 85%, implying a more balanced citation distribution across their publications. The small confidence intervals for these values reinforce the accuracy of our diversity calculations. Fig. 15(b) presents a box-plot of the total citations for each laureate, maintaining the same order of laureates as in Fig. 15(a) for easy comparison. Although the total citation ranges vary among the laureates, they do not provide significant insights into citation distribution. Instead, our diversity values effectively illustrate the extent of citation inequality among these economics laureates.
@@ IV.3.6 The Nobel Prize winners in Physiology/Medicine
In Fig. 16, we show the diversity values and total citation ranges for 15 recent Nobel laureates in Physiology/Medicine. Fig. 16(a) categorizes them into three groups: five laureates with diversity below 85%, eight between 85%-90%, and two above 90%, indicating low citation inequality overall. The minimal confidence intervals confirm the reliability of these diversity values. In Fig. 17, we examine diversity values and citation ranges of earlier laureates, with most diversity values falling between 65%-87%, except for a laureate from 1904 who shows significant citation inequality. The broader confidence intervals for these early laureates suggest greater uncertainty in the data, unlike the more reliable and precise values seen in recent times.
@@
In summary, this comprehensive analysis across multiple disciplines— physics, chemistry, mathematics, computer science, economics, and physiology or medicine demonstrates the significance of citation diversity values as a more insightful metric than total or average citation counts. In physics, our analysis illustrate that diversity values have risen over time, indicating decreased citation inequality among Nobel laureates. In chemistry, despite an increase in average citation counts, diversity values have remained stable, underscoring their role in revealing citation distribution. Similar findings in mathematics reinforce the relevance of this measure. For Turing award winners in computer science, our calculated diversity measure effectively captures citation distribution, providing clearer insights into citation inequality. In economics, high diversity values denote low citation inequality, offering a comprehensive understanding of laureates’ citation patterns. In physiology or medicine, diversity values have increased in recent times, indicating a more equitable distribution of citations among publications. Since the awards in mathematics, computer science and economics began relatively later we cannot provide a comparative analysis, over time, of the citation diversity values in these 3 disciplines. Our analysis advocate for the adoption of more nuanced metrics in evaluating scholarly impact, fostering a fairer assessment of academic contributions across various scientific fields.
@@ IV.4 Citation diversity in the publication of prize winning male and female scientists
Table 13 provides detailed information on the number of male and female award-winning scientists across different scientific disciplines, along with the period of our analysis. Additionally, the table includes the average citation count per publication for all male and female award winners in different disciplines and our calculated diversity values for these scientists. It is noted that among 126 recent award winners across six disciplines, there were no female scientists in mathematics and computer science during the period under consideration. In a graphical representation, Fig. 18(a) reveals that although both male and female scientists exhibit high diversity (lower inequality) values, indicating low citation inequality, male scientists generally have higher diversity values than female scientists in physics, economics, and physiology/medicine (in physics and physiology/medicine the diversity values of male are very close). In chemistry, however, female award winners show a more even citation distribution than their male counterparts. Both male and female scientists in physics and chemistry have high diversity values (above 90%). Conversely, in economics and physiology/medicine, there is a significant difference in diversity values, with female scientists experiencing higher citation inequality. Additionally, Fig. 18(b) depicts the total citation range for male and female award winners in these four disciplines. Given the greater number of male scientists, their total citation range is higher. However, when examining the average citation count per scientist, female scientists in chemistry have a higher average, supporting the diversity value findings. While the average citation count and total citation range provide some insights, the diversity values more effectively illustrate citation distribution and inequality among male and female award winners in each discipline.
@@ V Conclusion
Our extensive study on citation analysis sheds light on various aspects of global citation inequality, offering a detailed understanding of citation patterns across different countries and academic disciplines. Key highlights of our work may be summarized as follows:
@@
  • Distribution pattern of citation counts: We examined the distribution of citation counts in top institutes across various countries, revealing the Pareto law nature of the upper end of the distribution with a breakdown at the lower end. We also showed how the Pareto law’s scaling exponent changes with the number of top institutes across the globe.
@@
  • Novel citation diversity measure: A new log-normal entropy (LNE) based diversity measure has been introduced to measure citation inequality, enhancing the accuracy and depth of our analysis. Previous research has extensively explored diversity measures across various fields, also employing different metrics to assess citation distributions. However, this study marks the first instance of using an entropy-based diversity measure specifically to quantify citation distribution. We have utilized this innovative metric to effectively measure citation inequality, enhancing our understanding of the disparities in citation patterns.
@@
  • Institutional citation diversity measure across the world: We calculated citation diversity measures with confidence intervals, grouping countries based on these measures in top institutes (10, 20, 50) worldwide. This revealed that many small countries share groups with large economic powers, and groupings shift with the number of institutes considered. Box-plots were utilized to study the total number of citations, suggesting the emergence of subgroups based on citation counts.
@@
  • Discipline-wise citation diversity: We further calculated citation diversity along with total citation counts (box-plots) of award winning scientists in six disciplines (21 scientists from each discipline), uncovering the importance of measuring citation diversity of award winners across disciplines. Time evolution of the citation diversity across disciplines over the century was also studied in three main disciplines.
@@
  • Citation diversity of publications of award winning individual scientists: Citation diversity measures were analyzed for publications by award winning scientists in six disciplines (from 2000-2023 with publicly available data of 15 scientists from each disciplines), showing significant variation across fields. This was also extended to individual award winners from 1901-1920 in three principal disciplines. Also the time evolution of author wise diversity measures in three disciplines provided insights into how citation patterns change over time.
@@
  • Gender-based study in citation diversity: A gender-based analysis of citation diversity for male and female scientists in four disciplines highlighted areas of both low and high inequality, with a notable absence of female award winners in two disciplines during our considered period of analysis from 2007 to 2023.
@@
This extensive study, based on the data of the top institutes or highly acclaimed elite researchers, underscores the complexity and diversity of citation practices across the scientific landscape, offering a detailed examination that considers multiple dimensions and perspectives. Our findings suggest that the new citation diversity measure serves as a vital metric for assessing citation inequality, providing exceptional insights that citation counts alone cannot achieve. As a future research project, to portray the citation diversity analysis of the entire scientific community, one may incorporate the data from a larger and more diverse group of scientists, beyond just the elite group of top award winners. Additionally, the investigation of the lower end of citation distributions, either in isolation or in conjunction with other distributional models that adequately fit the overall data, is another important open research question for future research works. This could provide deeper insights into the factors influencing lower citation counts and shed light on the dynamics governing citation inequalities. Through this, we aim to foster a more clearer understanding of scientific contributions globally.

### v2

@@ IV.3.2 Nobel Prize winners in Chemistry
From 1901 to 2023 the Nobel prize in chemistry has been bestowed on a total of 192 individuals. For our analysis, we have picked up the citation data of 30 Nobel prize winners in chemistry, 15 during 1901-1920 and 15 between 2018-2023. Table 8 represents the calculated diversity values and average citation counts of all the listed Nobel prize winners in chemistry. In Fig. 11, we see the citation diversity of 15 recent Nobel laureates in chemistry, with diversity values ranging from 60% to 90%. Two distinct groups emerge: one containing four laureates with citation diversity between 60%-70%, indicating lower citation diversity, and another between 80%-90%, reflecting a more balanced citation distribution. The minimal confidence intervals confirm the reliability of these values. Fig. 11(b) further shows that despite similar average citation counts, diversity values differ significantly, revealing more insightful patterns in citation distribution. Meanwhile, Fig. 12 shows earlier laureates’ citation diversity ranging from 65% to 90%, though with larger confidence intervals, indicating greater variability.
@@ IV.3.3 Abel prize winners in Mathematics
The Abel prize is awarded annually (2003 onwards) to one or more outstanding mathematicians and is widely considered the Nobel prize of mathematics. Here we consider the data for 15 Able prize winners during 2012-2023. Table 9 presents the average citation counts and diversity values for each Abel prize winner considered. In Fig. 13, we observe the citation diversity and box-plot for the citation counts of each Abel prize winner. These citation diversity values of each scientist range from 70% to 90%, indicating moderate citation diversity. The confidence intervals for most diversity values are not high, though a few have slightly higher confidence intervals. This suggests that most calculated diversity values are reliable, with some variability due to the use of publicly available data.
@@ IV.3.4 Turing Award winners in Computer Science
In Fig. 14, we present an overview of the diversity values and citation ranges for each Turing award winner in computer science. scientists Notably, the diversity value of one of the 2015 Turing award winners is significantly lower compared to the others, indicating an uneven citation distribution for that individual scientist. Conversely, the diversity values of other computer scientists range between 55% and 85%, with a distinct division at 70%. Below this threshold, there are 7 Turing award winners, and above it, there are also 7 winners. This suggests that the citation diversity is lower for the 7 scientists (below 70%) compared to those above it. In this analysis, the confidence interval is minimal, except for two cases. For better understanding, we maintain the same order of the award winners on the x-axis in Fig. 14(b) as in Fig. 14(a), where they are arranged in ascending order of diversity values. However, no meaningful pattern is observed from the total citation range in Fig. 14(b) but there is a clear pattern in diversity values of each scientists shown in Fig. 14(a).
@@ IV.3.5 Nobel prize winners in Economics
The first Nobel Memorial Prize in economic sciences was awarded in 1969. As of 2023, 55 prizes in economic sciences have been given to 93 individuals. We considered the publication data of 15 distinguished economists between 2017-2023. Table 11 presents the diversity values and average citation counts for these 15 individual Nobel laureates in economics.
@@
Fig. 15(a) displays the citation diversity values for each laureate, showing that two recent laureates have very low diversity. In contrast, citation diversity of the other laureates ranging from 70% to 85%, implying a more balanced citation distribution across their publications. The small confidence intervals for these values reinforce the accuracy of our diversity calculations. Fig. 15(b) presents a box-plot of the total citations for each laureate, maintaining the same order of laureates as in Fig. 15(a) for easy comparison. Although the total citation counts vary among the laureates, it does not provide any significant insights into citation distribution. Instead, our diversity values effectively illustrate the extent of citation counts among these economics laureates.
@@ IV.3.6 The Nobel Prize winners in Physiology/Medicine
In Fig. 16, we show the diversity values and total citation ranges for 15 recent Nobel laureates in Physiology/Medicine. Fig. 16(a) categorizes them into three groups: five laureates with diversity below 85%, eight between 85%-90%, and two above 90%, indicating high citation diversity overall. The minimal confidence intervals confirm the reliability of these diversity values. In Fig. 17, we examine diversity values and citation ranges of earlier laureates, with most diversity values falling between 65%-87%, except for a laureate from 1904 who shows significant low citation diversity. The broader confidence intervals for these early laureates suggest greater uncertainty in the data, unlike the more reliable and precise values seen in recent times.
@@
In summary, this comprehensive analysis across multiple disciplines— physics, chemistry, mathematics, computer science, economics, and physiology or medicine demonstrates the significance of citation diversity values as a more insightful metric than total or average citation counts. In physics, our analysis illustrate that diversity values have risen over time among Nobel laureates. In chemistry, despite an increase in average citation counts, diversity values have remained stable, underscoring their role in revealing citation distribution. Similar findings in mathematics reinforce the relevance of this measure. For Turing award winners in computer science, our calculated diversity measure effectively captures citation distribution. In economics, high diversity values offering a comprehensive understanding of laureates’ citation patterns. In physiology or medicine, diversity values have increased in recent times, indicating a more equitable distribution of citations among publications. Since the awards in mathematics, computer science and economics began relatively later we cannot provide a comparative analysis, over time, of the citation diversity values in these 3 disciplines. Our analysis advocate the adoption of more nuanced metrics in evaluating scholarly impact, fostering a fairer assessment of academic contributions across various scientific fields.
@@ IV.4 Citation diversity in the publication of prize winning male and female scientists
Table 13 provides detailed information on the number of male and female award-winning scientists across different scientific disciplines, along with the period of our analysis. Additionally, the table includes the average citation count per publication for all male and female award winners in different disciplines and our calculated diversity values for these scientists. It is noted that among 126 recent award winners across six disciplines, there were no female scientists in mathematics and computer science during the period under consideration. In a graphical representation, Fig. 18(a) reveals that although both male and female scientists exhibit high diversity values, male scientists generally have higher diversity values than female scientists in physics, economics, and physiology/medicine (in physics and physiology/medicine the diversity values of male are very close). In chemistry, however, female award winners show a more even citation distribution than their male counterparts. Both male and female scientists in physics and chemistry have high diversity values (above 90%). Conversely, in economics and physiology/medicine, there is a significant difference in diversity values with female scientists. Additionally, Fig. 18(b) depicts the total citation for male and female award winners in these four disciplines. Given the greater number of male scientists, their total citation range is higher. However, when examining the average citation count per scientist, female scientists in chemistry have a higher average, supporting the diversity value findings. While the average citation count and total citation range provide some insights, the diversity values more effectively illustrate citation distribution among male and female award winners in each discipline.
@@ V Conclusion
Our extensive study on citation analysis sheds light on various aspects of global citation diversity, offering a detailed understanding of citation patterns across different countries and academic disciplines. Key highlights of our work may be summarized as follows:
@@
  • Distribution pattern of citation counts: We have examined the distribution of citation counts among top institutes across various countries, revealing the Pareto law nature of the upper end of the distribution with a breakdown at the lower end. It has also been showed how the Pareto law’s scaling exponent changes with the number of institutes considered across the globe.
@@
  • Novel citation diversity measure: A novel log-normal entropy (LNE) has been used to measure citation diversity in our analysis. Previous researches have extensively explored diversity measures across various fields, employing different metrics to assess citation distributions. However, this study marks the first instance of using an entropy-based diversity measure specifically to quantify citation distribution. We have utilized this innovative metric to effectively measure citation diversity, enhancing our understanding of the disparities in citation patterns.
@@
  • Institutional citation diversity measure across the world: We calculated citation diversity measures with confidence intervals, grouping countries based on these measures with respect to top few (10, 20, or 50) institutes worldwide. This revealed that many small countries share groups with large economic powers, and these groupings shift with the number of institutes considered. Box-plots have been utilized to study the total number of citations, suggesting the emergence of subgroups based on citation counts.
@@
  • Discipline-wise citation diversity: We further calculated citation diversity along with total citation counts of award winning scientists in six disciplines (21 scientists from each discipline), physics, chemistry, mathematics, computer science, economics and physiology/medicine, uncovering the importance of measuring citation diversity of award winners across disciplines. Time evolution of the citation diversity across disciplines over the century has also been studied in three main disciplines (physics, chemistry and physiology/medicine).
@@
  • Citation diversity of publications of award winning individual scientists: Citation diversity measures have analyzed for publications by award winning scientists in six disciplines (from 2000-2023 with publicly available data of 15 scientists from each disciplines, physics, chemistry, mathematics, computer science, economics and physiology/medicine), showing significant variation across fields. This has also been extended to individual award winners from 1901-1920 in three principal disciplines. The time evolution of author-wise diversity measures in three disciplines (physics, chemistry and physiology/medicine) provides insights into how citation patterns change over time.
@@
  • Gender-based study in citation diversity: Finally, a gender-based analysis of citation diversity, during the period 2007-2023, has been done for male and female scientists in four disciplines (physics, chemistry, economics and physiology/medicine). The absence of female award winners in two disciplines (mathematics and computer science) has been noted in the considered period of our analyses.
@@
This extensive study, based on the data of the top institutes or highly acclaimed elite researchers, underscores the complexity and diversity of citation practices across scientific landscapes, offering a detailed examination from multiple dimensions and perspectives. Our findings suggest that the new measure of citation diversity serves as a vital metric to assess the unevenness of the citation distribution, providing exceptional insights that citation counts alone cannot achieve. The diversity measure, D (ranging from 0 to 100) quantifies the uniformity in the distributions of the citation patterns. Higher values of D indicate more balanced distributions, while lower values suggest concentration among a few institutions, or a few research articles as the case may be. As a future research project, to portray such citation diversity analysis of the entire scientific community, one may incorporate the data from a larger and more diverse group of scientists, beyond just the elite group of top award winners. Further, our findings could be compared with established inequality indices to gain deeper insights into the structural unevenness within the scientific community. Additionally, investigating the lower end of the citation distributions, either in isolation or in conjunction with other distributional models that adequately fit the overall data, is another important open research question for future work. This could provide deeper insight into the factors driving lower citation counts and shed light on the dynamics that govern citation disparities.
Acknowledgement
We have submitted the same manuscript on arXiv as a pre-print version [8]. The authors appreciate constructive suggestions of three anonymous reviewers.